OpenStudy (anonymous):

Solve the equation for x if ax + bx = c x = c/(a + b) x = (a + b)/c x = c - a - b x = c/(ab)

OpenStudy (anonymous):

which awnser is it please tell me

OpenStudy (nirmalnema):

x = c/(a + b)

OpenStudy (anonymous):

Well let's isolate the x.

OpenStudy (anonymous):

Don't just give answers please +nirmalnema it is against the Code of Conduct, try to lead them to solutions so they understand what they are doing - and also so you dont get banned from OpenStudy

OpenStudy (nirmalnema):

ok

OpenStudy (anonymous):

Ok so we have: ax + bx = c And we want to solve for x So what we must do is isolate x to one side of the equation. We do this by first factoring out x from ax + bx like so: x(a+b) = c

OpenStudy (anonymous):

sorry but its x=c/(a+b)

OpenStudy (anonymous):

Now we must divide both sides by (a+b) to get x by itself.